OpenStudy (anonymous):

Simplify: (2+3i)+(2-3i) A. 5 B. 4+6i C. 4 D. 11

OpenStudy (anonymous):

Just add and subtract like normally, dropping the parentheses we get, 2+2 = 4, 3i-3i = 0, so therefore your answer is 4.

OpenStudy (raden):

by using difference square formula : (a+b)(a-b) = a^2 - b^2 so, (2+3i)+(2-3i) = 2^2 - (3i)^2 = 4 - 9i^2 = 4 - 9*(-1) = 11 since that i = sqrt(-1) then i^2 = -1

OpenStudy (anonymous):

I don't think i is treated here as a imaginary number, just a variable. As you can see from the given options :p

OpenStudy (anonymous):

Oh wait hold on, you're right o_O, it should be 11.

OpenStudy (raden):

no, im wrong... damn, i saw mistake in the question. it is (2+3i)+(2-3i) not (2+3i)(2-3i) haha :v

OpenStudy (anonymous):

Ahaha alright, I started thinking that soon as you posted that as well >_>

OpenStudy (anonymous):

bleh lol

OpenStudy (anonymous):

Thanks for checking though!